OpenStudy (anonymous):

What is the equation of a horizontal line passing through the point (-3, 8)? b. y = 8 c. x = 5 d. y = -11

OpenStudy (anonymous):

@phi How do I do this one

OpenStudy (anonymous):

y = 8

OpenStudy (anonymous):

Yeah but I wanna know how you got there

OpenStudy (anonymous):

@sourwing @phi

OpenStudy (phi):

a horizontal line goes sideways (not up or down) y measures up or down if a horizontal line goes through (-3, 8) its y value is 8 (2nd number) because it is horizontal its y value never changes. y=8 if you think of this in slope intercept form y= mx + b b is 8 and m is 0 (slope is 0) y = 0*x + 8 or just y= 8

jimthompson5910 (jim_thompson5910):

all horizontal lines are of the form y = k where k is some number to make sure that y = k goes through (-3,8), k must be 8 since this is the y coordinate of the given point So that's why the answer is y = 8

jimthompson5910 (jim_thompson5910):

basically, every point on y = 8 is a point of the form (x,8) x is any number and the y coordinate is fixed to be 8

OpenStudy (phi):

The vertical line is the same way, but in the form x= number

OpenStudy (anonymous):

Oh....well I feel extremely stupid now

OpenStudy (phi):

people are not born knowing this